Ivanhoe Company has a net profit margin of 11.8 percent, debt ratio of 53 percent, total assets of $4,048,200, and sales of $7,432,500. If the company has a dividend payout ratio of 75 percent, what is its sustainable growth rate?

Sustainable growth rate %
Answer #1

Equity =Total Assets*(1-Debt Ratio) =4048200*(1-53%) =1902654
Net Profit =Profit Margin*Sales =11.8%*7432500 =877035
ROE =Net Profit Margin/Equity =877035/1902654 =46.0953%
b =1- Dividend payout ratio = 1-75% =25%
Sustainable Growth Rate =ROE*b/(1-ROE*b) =46.0953%*25%/(1-46.0953%*25%) =13.02%